OpenStudy (anonymous):

In a a parallelogram DEFG, DH=y+3, HF=2x+1, GH=2y-1, and HE=x+4 what are values of y and x

OpenStudy (anonymous):

I'll give medal!!

pooja195 (pooja195):

Hi Welcome To Openstudy :-) Any ideas on where to start?

OpenStudy (anonymous):

No:(

pooja195 (pooja195):

It's alright ^_^ Ok so we can say H is the center of both parallelogram diagonals

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

ok so we have this: DH = HF \[\huge~\rm~y+3=2x+1\] isolate y

OpenStudy (anonymous):

X=y/2+1?

pooja195 (pooja195):

Yes we can also write it as x=1/2y+1 <---leave it as this for right now Next we know the GH=HE So... \[\huge~\rm~ 2y-1=x-4\] isolate the y

OpenStudy (anonymous):

Y=x/2-3/2

OpenStudy (anonymous):

I think x =3 and y =4

OpenStudy (anonymous):

Wait no

pooja195 (pooja195):

sorry isolate the 2nd equation for x

pooja195 (pooja195):

not y

OpenStudy (anonymous):

X=2y+3

pooja195 (pooja195):

Good now we have 2 equations x=1/2y+1 X=2y+3 Now we use substitution form here think you can do it? :)

pooja195 (pooja195):

Wait what are your answer choices?

OpenStudy (anonymous):

X=14 and y=11, x=3 and y = 4, and x=7 and y. =4, x=1 and y =. 11

pooja195 (pooja195):

This? 

OpenStudy (anonymous):

Not the same question

pooja195 (pooja195):

Ok anyways... Good now we have 2 equations x=1/2y+1 X=2y+3 Now we use substitution form here think you can do it? :)

pooja195 (pooja195):

We isolated wrong y+3=2x+1 <---this is what we are supposed to have

pooja195 (pooja195):

x=2y-5 and x=(y+2)/2 would be our 2 equations

OpenStudy (anonymous):

Oh yeah

pooja195 (pooja195):

Now try solving :)

OpenStudy (anonymous):

X = 3 and y=4

pooja195 (pooja195):

Can you show me how you got that?

OpenStudy (anonymous):

I substituted, 3=2(4)-5 and 3=4+2/2

imqwerty (imqwerty):

nice :)

pooja195 (pooja195):

Perfect :)

pooja195 (pooja195):

Good job ^_^ \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

Thanks ! Happy holidays

pooja195 (pooja195):

Same to you! :D and yw ^_^